OpenStudy (anonymous):

Find f(g(x) when f(x) is 3/x+1 and g(x) is 2/x I plugged in the numbers and got : 3/(2/x)-1

OpenStudy (jhannybean):

f(g(x)) = [3/(2/x)] +1 \(\checkmark\) You just have to simplify it. \[f(g(x)) = \dfrac{3}{\frac{2}{x}}-1\]Remember: b/(a/c) = bc/a With this concept, we can simplify the first part of the function. \[f(g(x)) = \frac{3x}{2}-1\] Or \[f(g(x)) = \frac{3x-2}{2}\] Depends how you'd prefer writing it, and the second one is just combining both terms into one form.